OpenStudy (anonymous):

What is the simplified form of ((2/x)+(3/y))/((2/x)-(3/y))

OpenStudy (anonymous):

That IS the simplified form....

OpenStudy (anonymous):

\[\frac{(\frac{2}{x} + \frac{3}{y})}{(\frac{2}{x} - \frac{3}{y})}\]

OpenStudy (anonymous):

Oh, I read it wrong.....

OpenStudy (anonymous):

\[\frac{(\frac{2}{x} + \frac{3}{y})}{1} \times \frac{1}{(\frac{2}{x} - \frac{3}{y})}\] Just work from the inside out

OpenStudy (anonymous):

multiply top and bottom by \(xy\) cancelling merrily as you got you will be done in one step

OpenStudy (anonymous):

\[\frac{(\frac{2}{x} + \frac{3}{y})}{(\frac{2}{x} - \frac{3}{y})}\times \frac{xy}{xy}\]

OpenStudy (anonymous):

Or the very long way: